# Vipera: Blending Visual and LLM-Driven Guidance for Systematic Auditing of Text-to-Image Generative AI

### v1

@@ 5.1. Study Design and Methodology
Participants. We recruited 24 participants (P1-P24, 14 males and 10 females). Among them, 20 were general student auditors recruited from our institution (4 undergraduates, 7 Master’s, and 9 Ph.D students) who reported an average familiarity with AI auditing (M=3.04/5, SD=1.16) and use T2I models with an average frequency (M=2.92/5, SD=0.929). The remaining 4 participants were expert auditors recruited from both academic and industry who had rich experience and knowledge in AI auditing. The ages of the participants ranged from 20-29, with an average of 23.9.
@@
Figure 6. The additional Criteria View added to Systems A and C. Users are allowed to add auditing criteria, and the results will be visualized as bar charts in the respective cards. Each criterium card includes three buttons for editing the criterium, relabeling the images, and bookmarking the criterium, respectively.:
@@
Tasks. Participants were asked to audit the Stable Diffusion XL model. They were assigned one of four initial prompts designed to elicit a range of potential biases and quality issues: (a) A couple on their wedding day, (b) A family having a picnic in the park, (c) Worldwide athletes in the Olympic Games, or (d) An award-winning chef preparing a gourmet meal. These prompts are able to cover diverse topics and image features, such as individuals (a, d), and crowd (b, c), country (c), occupation bias (a, c), cultural issues and potential bias (c, d). Participants were instructed to explore the model’s behavior freely from this starting point.
@@
Procedure. Each session lasted approximately 90 minutes. After providing informed consent, participants completed a demographic questionnaire. They were then equally divided into two groups (Group I and Group II), with the distribution of general and expert auditors kept consistent across both groups. The study employed a mixed design where participants in Group I used systems A, B, and D, while those in Group II used systems A, C, and D. Before using each system, they were provided a tutorial on the system, followed by a warm-up session of up to 5 minutes to familiar themselves with the interface. Subsequently, they were given 15 minutes to perform an auditing task, with the goal of acquiring as many insights as possible. To balance between controlled comparison and free exploration, all participants began with their assigned prompt but were free to write subsequent prompts on the same topic. Each starting prompt was assigned to an equal number of participants. They were also instructed to note down their insights in phrases or short sentences as soon as they get them to eventually form an auditing report after using each system. We employed the think-aloud protocol (Lewis and Mack, 1982) to capture their reasoning process. After completing the tasks with all three systems, participants filled out a questionnaire that included NASA-TLX scales for each system and several 7-point Likert scales to rate the usefulness of system components. The study concluded with a semi-structured interview to gather qualitative feedback. Participants received a \$10 gift card as the compensation for their time. All sessions were both screen- and audio-recorded. The study proposal has been approved by the IRB committee of our institution.
@@
Data collected and analysis. We collected multiple forms of data for each participant: three auditing reports (one for each system used), pre- and post-study questionnaires, system interaction logs, and full session recordings. Our analysis approach was twofold. Quantitative data from the questionnaire responses and system interaction logs were analyzed using statistical methods to assess performance and usability. Qualitative data from the auditing reports and transcripts from the think-aloud protocol and semi-structured interviews, were analyzed using thematic analysis (Braun and Clarke, 2006) to identify user strategies, reasoning patterns, and key feedback themes.
@@ 6. Results
In this section, we share the results from our controlled experiments. In particular, analysis of the NASA-TLX questionnaire responses revealed that participants reported lower workload and improved performance when using Vipera—especially with the AI auditing support—compared to the baseline. Auditing logs show that Vipera helped participants create more auditing criteria through interaction with AI support and visual guidance. Our interviews with practitioners also revealed nuanced trade-offs between scaffolding systematic auditing while managing cognitive demand, personalizing recommendations, and leveraging resources tailored to individual auditors.
@@ 6.1.1. Questionnaires
The detailed distribution of participants’ rating on NASA-TLX scales is illustrated in Figure 7. Overall, participants reported less workload in all NASA-TLX dimensions and better performance on average when using Vipera (D) compared with the baseline (A). The difference was significant in the Mental Demand (t=-2.055, p=0.0257), Physical Demand (t=-2.061, p=0.0254), and Performance (t=2.685, p=0.0066) dimensions. These results show that Vipera is generally effective in reducing users’ perceived workload and improve the auditing performance.
For participants in Group I, we observed a general pattern of reduced Physical Demand and Temporal Demand and increased Performance across the sequence from A to D as shown in the Group I panel. Specifically, Performance improved significantly from A to B (t = 3.023, p = 0.0058), and from B to D the reported increase in Performance was also significant (t = 1.876, p = 0.0437). Note that the latter p-value is modest and multiple comparisons were performed; if corrections for multiple testing are applied the significance of some comparisons may change. It is worth noting that the Mental Demand in A and B has the same median with the average of B slightly higher, and meanwhile the median of both of them are higher than D. Overall, these results suggest that the scene graph alone may increase mental demand, while subsequent addition of the AI auditing support (D) reduces workload and improves performance, but order and multiple-comparison concerns warrant cautious interpretation.
Figure 7. Participants’ ratings on the NASA-TLX scale. The first subplot shows a rating comparison ratings of all participants on the Baseline (A) and Vipera (D), while the other two subplots shows detailed rating comparison for the three systems used by participants in each group. The white marks on the boxes indicate the medians. :
@@
Meanwhile, for participants in Group II, System C generally shows a lower workload than System A across multiple dimensions, including Mental Demand, Physical Demand, Temporal Demand, and Effort. Figure 7 also indicates that Performance and Frustration across A, C, and D are broadly comparable with overlapping distributions. Statistical comparisons show that participants reported significantly lower Mental Demand (t = -5.063, p = 0.0002), Physical Demand (t = -1.959, p = 0.038), Temporal Demand (t = -2.077, p = 0.031), and Effort (t = -2.803, p = 0.0086) for System C compared with System A. They also reported significantly higher Temporal Demand for System D compared with C (t = 2.569, p = 0.0130). These results indicate that the AI auditing support can effectively reduce users’ mental workload and effort when added alone, while the scene graph may introduce additional temporal demand when added alongside the AI auditing support. Most participants attributed this extra temporal demand to the significant amount of time spent on exploring promising auditing criteria when the scene graph is present. A few participants also mentioned the fact that it is naturally difficult to find useful criteria and get insights in systems used at a later time when they have already explored the model extensively in the previous systems.
@@
In addition, participants generally perceived all components within Vipera as helpful, with an average rating of 3.50 (SD=0.885) and 3.67 (SD=0.868) on the scene graph and AI auditing support, respectively. They also appreciated the automatic labeling of images (M=3.375, SD=0.970) and note-taking features (M=3.71, SD=1.083).
@@ 6.1.2. Auditing logs
One interesting discovery is that the introduction of AI auditing support (System A vs. C; System B vs. D) consistently leads to an increase in the number of prompts, images, and criteria, as well as a consistent decrease in bookmarked charts or images. This shows that auditing support powered by AI is beneficial and inspirational for most participants despite the extra effort in exploration. By contrast, the effect brought by the introduction of the scene graph (System A vs. B; System C vs. D) was more diverse and random, varying from person to person.
Figure 8. The number of prompts, images, used criteria, and bookmarked charts or images in each system used by the participants. Error bars indicate standard error of the mean.:
6.1.3. Auditing reports
It’s worth noting that although we clarified in the system tutorial that the task was to audit the T2I model specifically, a few participants still misunderstood the goal and commented on other system components, such as the scene graph and the AI labeler. Relatedly, while many participants noticed the errors of the AI labeler, only about half of them tried to correct the errors or asked for relabeling with modified expressions, with many participants ending up trusting the wrong results. This pattern extends prior work on aligning user intents with LLM-based judges (Shankar et al., 2024; Pan et al., 2024), showing that even when users are prompted to evaluate a specific component, failures in automated judges can go unchallenged, so such tools should be treated and deployed with caution.
@@ 6.2. Qualitative Findings from Interviews
We received divided opinions on the effect of the scene graph. Participants in favor of this visual guidance suggested that it provided them with a general overview of the images, served as a good source of inspiration, and encouraged systematic thinking and attention to detail, ultimately making them feel more productive and confident. For instance, P1 noted, “I was used to focusing on the macro scenario in the image and treating it as a whole, but the scene graph made me realize that images could be decomposed into individual elements.” P21 similarly mentioned that the scene graph gave him “a systematic understanding of the images’ compositions”. Moreover, we observed that some participants acquired inspirations from interacting with the scene graph. P3 mentioned that apart from the tree structure and the node contents, the results of open labeling (i.e., without predefined candidate values) could also be inspiring, reminding him of “corner cases”. However, others found the scene graph’s complexity to be a barrier, citing a steep learning curve and information overload. This structured approach also had a psychological cost for some. P06 and P09, for instance, felt pressured by the graph, stating it limited the “free-form authoring” he preferred and undermined his sense of agency. Additionally, several participants expressed neutral views. P21 and P08 argued that the usefulness of the scene graph was most prominent when the prompts or images were abstract (i.e., with unclear scenes or in an artistic style), which were easier to assess using the scene graph compared with direct inspection. P21 further indicated that the scene graph’s usefulness was content-dependent, asserting that nodes related to human characters or biases were more beneficial than those focusing on minor details.
@@ 7. Limitations
Our study has several limitations. First, we have identified several threats to the validity of our study. For instance, our controlled experiment, conducted in a lab-like setting with a fixed duration, may not fully capture the open-ended and long-term nature of real-world auditing tasks. Meanwhile, our participant pool consisted primarily of student auditors, whose workflows may differ from those of professional red teamers or compliance officers. The generalizability of our findings could be strengthened by future longitudinal deployments with a more diverse range of experts and domain-specific prompts. Furthermore, we used in our analysis bookmarks as a proxy for insights, a metric that future work could refine. The fixed order of system presentation could also introduce potential learning or fatigue effects, as the difficulty of finding new insights naturally increases over the course of a session. Finally, while Vipera can be extended to audit multiple T2I models in parallel, doing so likely requires further work to adapt guidance and reconcile differences across models.
@@ 8. Discussion
Based on our results, in this section we discuss opportunities to better support the design, development, and evaluation of AI auditing tools that enable more systematic auditing. In particular, we highlight how visual guidance can scaffold auditors before, during, and after their work by making complex results more interpretable and actionable. We also examine how AI-driven inspirations and guidance can be designed to better adapt to users’ goals, drawing on our empirical findings. In addition, we shed light on best practices for integrating different forms of guidance to support human–AI collaboration in auditing. Finally, we discuss the potential to embed tools like Vipera into on-the-ground responsible AI workflows in industry settings and beyond. Together, these opportunities outline design directions for building future auditing tools capable of scaling to the complexity and diversity of generative AI systems, while enhancing and complementing auditors’ capacity of sensemaking AI outputs at scale.

### v2

@@ 5.1. Study Design and Methodology
Participants. We recruited 24 participants (P1-P24, 14 males and 10 females). Among them, 20 were general student auditors recruited from our institution (4 undergraduates, 7 Master’s, and 9 Ph.D students) who reported an average familiarity with AI auditing (M=3.04/5, SD=1.16) and use T2I models with an average frequency (M=2.92/5, SD=0.929). The remaining 4 participants were expert auditors recruited from both academia and industry who had rich experience and knowledge in AI auditing. The ages of the participants ranged from 20-29, with an average of 23.9.
@@
Figure 6. The additional Criteria View added to Systems A and C. Users are allowed to add auditing criteria, and the results will be visualized as bar charts in the respective cards. Each criterion card includes three buttons for editing the criterion, relabeling the images, and bookmarking the criterion, respectively.
In this view, there is a view title ”Criteria” and an ”Add” button next to it for users to add an auditing criterion. There are currently two cards in this view, one for each criterium: a ”gender” card embedding a bar chart showing 15 male (blue) and 15 female (orange) in parallel, and a ”style” card embedding a bar chart showing 30 images under the ”cinematic” label (15 blue and 15 orange ones, visualized as a stacked bar chart. On the top right of each card, there are three buttons: edit, relabel, and bookmark.
@@
Tasks. Participants were asked to audit the Stable Diffusion XL model. They were assigned one of four initial prompts designed to elicit a range of potential biases and quality issues: (a) A couple on their wedding day, (b) A family having a picnic in the park, (c) Worldwide athletes in the Olympic Games, or (d) An award-winning chef preparing a gourmet meal. These prompts are able to cover diverse topics and image features, such as individuals (a, d), crowd (b, c), country (c), occupation bias (a, c), cultural issues and potential bias (c, d). Participants were instructed to explore the model’s behavior freely from this starting point.
Figure 7. Participants’ ratings on the NASA-TLX scale. The first subplot shows a rating comparison of all participants on the Baseline (A) and Vipera (D), while the other two subplots show a detailed rating comparison for the three systems used by participants in each group. The white marks on the boxes indicate the medians.
In the first subplot, the box plot shows that participants perceived less workload and better performance. The second subplot shows that participants perceived a decreasing level of workload from A, B, to D in Mental, Physical, Time, Effort and an increasing level of performance. For Frustration, the median of A is similar to D, and both are smaller than the median of B. The third subplot shows that participants perceived similar workloads in Time and Frustration and similar Performance. The Mental, Physical, and Effort workload of C and D is smaller than A, and there is no significant difference between C and D.
Procedure. Each session lasted approximately 90 minutes. After providing informed consent, participants completed a demographic questionnaire. They were then equally divided into two groups (Group I and Group II), with the distribution of general and expert auditors kept consistent across both groups. The study employed a mixed design where participants in Group I used systems A, B, and D, while those in Group II used systems A, C, and D. This progressive design allows us to observe how their auditing process and performance evolve as new capabilities are introduced. While a counterbalanced design in system order may mitigate the learning effect, it may introduce a more problematic contrast effect where participants evaluate the simpler system based on its missing features, rather than its intrinsic utility.
Before using each system, they were provided a tutorial on the system, followed by a warm-up session of up to 5 minutes to familiarize themselves with the interface. Subsequently, they were given 15 minutes to perform an auditing task, with the goal of acquiring as many insights as possible. To balance between controlled comparison and free exploration, all participants began with their assigned prompt but were free to write subsequent prompts on the same topic. Each starting prompt was assigned to an equal number of participants. They were allowed to use the same prompts across systems so we could more directly attribute differences in user strategy, efficiency, and audit outcomes to the system itself, clarifying each component’s impact on the user’s workflow for a consistent task.
@@
Participants were also instructed to note down their insights in phrases or short sentences as soon as they got them to eventually form an auditing report after using each system. We employed the think-aloud protocol (Lewis and Mack, 1982) to capture their reasoning process. After completing the tasks with all three systems, participants filled out a questionnaire that included NASA-TLX scales for each system and several 7-point Likert scales to rate the usefulness of system components. The study concluded with a semi-structured interview to gather qualitative feedback. Participants received a \$10 gift card as compensation for their time. All sessions were both screen- and audio-recorded. The study proposal has been approved by the IRB committee of our institution.
@@
Data collected and analysis. We collected multiple forms of data for each participant: three auditing reports (one for each system used), pre- and post-study questionnaires, system interaction logs, and full session recordings. Our analysis approach was twofold. Quantitative data from the questionnaire responses and system interaction logs were analyzed using statistical methods to assess performance and usability. Qualitative data from the auditing reports and transcripts from the think-aloud protocol and semi-structured interviews were analyzed using thematic analysis (Braun and Clarke, 2006) to identify user strategies, reasoning patterns, and key feedback themes.
@@ 6. Results
In this section, we share the results from our controlled experiments. In particular, analysis of the NASA-TLX questionnaire responses revealed that participants reported lower workload and improved performance when using Vipera—especially with the AI auditing support—compared to the baseline. Auditing logs indicate that Vipera assisted participants in creating more auditing criteria through interaction with AI support and visual guidance. Our interviews with practitioners also revealed nuanced trade-offs between scaffolding systematic auditing while managing cognitive demand, personalizing recommendations, and leveraging resources tailored to individual auditors.
@@ 6.1. Mixed-method Analysis of Questionnaires and Audit Logs
Figure 8. The number of prompts, images, used criteria, and bookmarked charts or images in each system used by the participants. Error bars indicate standard error of the mean.
There are three subplots in this figure. The first subplot shows the number of prompts, images, used criteria, and bookmarked charts or images in each system used by the participants of Group I (i.e., the systems are A, B, and D). For each system on the x-axis, there are four bar charts indicating the number of prompts, charts, criteria, and images from left to right, each with an error bar. Similarly, the second subplot corresponds to participants in Group II (i.e., systems A, C, and D) with a similar layout, while the third subplot corresponds to all participants (i.e., systems A, B, C, and D).
@@ 6.1.1. Questionnaires
The detailed distribution of participants’ ratings on NASA-TLX scales is illustrated in Figure 7. We analyzed ratings using paired t-tests, with p-values ($p$) corrected for multiple comparisons using the Holm–Bonferroni method ($p^{\prime}$). Our analysis first compared the final Vipera system (D) against the baseline (A) for all participants, then examined the step-wise introduction of components within each group to understand their individual contributions.
Our primary finding is that Vipera (D) delivered a statistically significant improvement in participants’ self-rated Performance compared to the baseline (A) ($t(23)$ = 2.685, $p$ = 0.0066, $p^{\prime}$ = 0.0397). While participants also reported lower average workload with Vipera, particularly in Mental Demand ($t(23)$ = -2.055, $p$ = 0.0257, $p^{\prime}$ = 0.0514) and Physical Demand ($t(23)$ = -2.061, $p$ = 0.0254, $p^{\prime}$ = 0.0762), these reductions represented a positive trend but were not statistically significant after correction.
To understand the naunces of why a clear performance gain does not accompany a corresponding significant drop in overall workload, we analyzed the data from each experimental group. The results reveal that the two main components of Vipera have distinct and somewhat opposing effects on user workload.
For Group II, introducing the AI auditing support alone (System C) caused a dramatic and statistically robust reduction in Mental Demand ($t(11)$ = -5.063, $p$ = 0.0002, $p^{\prime}$ = 0.0022) compared to the baseline. This powerful effect, which remained highly significant even after correcting for 12 planned comparisons, establishes the AI component as the primary driver of workload reduction in Vipera. Other workload dimensions like Effort ($p$ = 0.0086, $p^{\prime}$ = 0.0515), Physical Demand ($p$ = 0.0380, $p^{\prime}$=0.0912), and Temporal Demand ($p$ = 0.0310, $p^{\prime}$ = 0.0931) also showed trends toward reduction.
In contrast, the experience of Group I participants, who received the scene graph first (System B), was more mixed. We observed trends towards improved Performance when adding the scene graph (A vs. B, $t(11)$ = 3.023, $p$ = 0.0058, $p^{\prime}$ = 0.0696) or subsequently adding the AI support (B vs. D, $t(11)$ = 1.876, $p$ = 0.0437, $p^{\prime}$ = 0.1748), although neither remained robust after correction. We hypothesize that this inconsistency may be due to varying degrees of benefit derived from the scene graph across participants. In fact, descriptive data shows that median Mental Demand was slightly higher for System B than for the baseline (A) or Vipera (D), suggesting the scene graph may introduce its own cognitive overhead. However, subsequent addition of the AI auditing support reduces workload and improves performance on average.
@@
To summarize, the results indicate that the AI auditing support can reduce users’ mental workload, while the scene graph can contribute to auditing performance despite potential addition to the temporal demand. Most participants attributed this extra temporal demand to the significant amount of time spent on exploring promising auditing criteria when the scene graph is present. A few participants also mentioned the fact that it is naturally difficult to find useful criteria and get insights in systems used at a later time when they have already explored the model extensively in the previous systems.
@@
Finally, participants generally perceived all components within Vipera as helpful, with an average rating of 3.50 (SD=0.885) and 3.67 (SD=0.868) on the scene graph and AI auditing support, respectively. They also appreciated the automatic labeling of images (M=3.375, SD=0.970) and note-taking features (M=3.71, SD=1.083), indicating the overall high usability of Vipera interface.
@@ 6.1.2. Auditing logs
One interesting discovery is that the introduction of AI auditing support (System A vs. C; System B vs. D) consistently leads to an increase in the number of prompts, images, and criteria, as well as a consistent decrease in the number of bookmarked charts or images. This shows that auditing support powered by AI is beneficial and inspirational for most participants despite the extra effort in exploration. By contrast, the effect brought by the introduction of the scene graph (System A vs. B; System C vs. D) was more diverse and random, varying from person to person.
6.1.3. Auditing reports
Figure 9. Example auditing reports from one participant, P21. Reports are partially truncated due to the length.
In the first report (System A), the participant used one prompt ”A family having a picnic in the park” to generate 10 images, and wrote the following general notes: ”There seems to be a diverse group of people, including Asian and White but not so diverse to include other races. It could also be attributed to the artistic style of the images, which make race hard to tell. Prefer to see more realistic images but the majority (7) is artistic. The model does very well on the background and food. They seem pretty high-quality. It performs poorly on the facial expression which is something very important. The vibe is pretty accurate - families seem to have a good time.” In the second report (System B), the participant used two prompts to generate 10 images each: ”A family having a picnic in the park” and ”Two families having picnic in the park”. The general notes was ”It still leans heavily on artistic style when I don’t specify the style in the prompt”. Furthermore, a bar chart ”foreground.picnic_blanket.color” was cited with 8 ”red and white” and 2 ”others”. The participant wrote, ”It’s interesting that most picnic blankets are red and white. I probably want a more diverse set of colors. Could be attributed to most of the training images having this color”. In the third report (System D), the participant used two prompts to generate 10 images and 5 images respectively: ”A family having a picnic in the park” and ”A family having picnic in the park surrounded by flowers”. The general notes was ”The generated images are all in the same perspective, lacking some diversity. All images are in similar lighting condition. It’s not a bad thing because soft and bright light is common for picnics…”. A bar chart ”foreground.overhead_view” was also cited showing that all images were not taken from overhead.
Generally, we found the multi-modal guidance employed in Vipera enhanced the insight diversity in the audit reports, as illustrated by an example case shown in Figure 9. For instance, P21 initially focused on assessing the image quality, vibe, and race diversity in System A. The criticism of the model’s tendency to generate artistic images remained in the report from System B, but he also mentioned the issue within the diversity of blanket colors, citing a chart from the scene graph. In addition, the LLM-driven guidance in System D further inspired him to find that all images were taken from the front, which could be improved for better diversity.
It’s worth noting that although we clarified in the system tutorial that the task was to audit the T2I model specifically, a few participants still misunderstood the goal and commented on other system components, such as the scene graph and the AI labeler. For instance, as highlighted in Figure 9, P21 included a sentence commenting on the labeling quality in the report for System A. Relatedly, while many participants noticed the errors of the AI labeler, only about half of them tried to correct the errors or asked for relabeling with modified expressions, with many participants ending up trusting the wrong results. This pattern extends prior work on aligning user intents with LLM-based judges (Shankar et al., 2024; Pan et al., 2024), showing that even when users are prompted to evaluate a specific component, failures in automated judges can go unchallenged, so such tools should be treated and deployed with caution.
@@ 6.2. Qualitative Findings from Interviews
We received divided opinions on the effect of the scene graph. Participants in favor of this visual guidance suggested that it provided them with a general overview of the images, served as a good source of inspiration, and encouraged systematic thinking and attention to detail, ultimately making them feel more productive and confident. For instance, P1 noted, “I was used to focusing on the macro scenario in the image and treating it as a whole, but the scene graph made me realize that images could be decomposed into individual elements.” P21 similarly mentioned that the scene graph gave him “a systematic understanding of the images’ compositions”. Moreover, we observed that some participants acquired inspiration from interacting with the scene graph. P3 mentioned that apart from the tree structure and the node contents, the results of open labeling (i.e., without predefined candidate values) could also be inspiring, reminding him of “corner cases”. However, others found the scene graph’s complexity to be a barrier, citing a steep learning curve and information overload. This structured approach also had a psychological cost for some. P06 and P09, for instance, felt pressured by the graph, stating it limited the “free-form authoring” he preferred and undermined his sense of agency. Additionally, several participants expressed neutral views. P21 and P08 argued that the usefulness of the scene graph was most prominent when the prompts or images were abstract (i.e., with unclear scenes or in an artistic style), which were easier to assess using the scene graph compared with direct inspection. P21 further indicated that the scene graph’s usefulness was content-dependent, asserting that nodes related to human characters or biases were more beneficial than those focusing on minor details.
@@ 7. Limitations
Our study has several limitations. First, we have identified several threats to the validity of our study. For instance, our controlled experiment, conducted in a lab-like setting with a fixed duration, may not fully capture the open-ended and long-term nature of real-world auditing tasks. Meanwhile, our participant pool consisted primarily of student auditors, whose workflows may differ from those of professional red teamers or compliance officers. The generalizability of our findings could be strengthened by future longitudinal deployments with a more diverse range of experts and domain-specific prompts. Furthermore, we used bookmarks in our analysis as a proxy for insights, a metric that future work could refine. The fixed order of system presentation could also introduce potential learning or fatigue effects, as the difficulty of finding new insights naturally increases over the course of a session. In addition, the LLMs used in the study could make mistakes that misled participants, which could influence user responses and auditing results. Finally, while Vipera can be extended to audit multiple T2I models in parallel, doing so likely requires further work to adapt guidance and reconcile differences across models.
@@ 8. Discussion
Based on our results, this section discusses opportunities to enhance the support for the design, development, and evaluation of AI auditing tools that facilitate more systematic auditing. In particular, we highlight how visual guidance can scaffold auditors before, during, and after their work by making complex results more interpretable and actionable. We also examine how AI-driven inspirations and guidance can be designed to better adapt to users’ goals, drawing on our empirical findings. Additionally, we provide insights into best practices for integrating various forms of guidance to support human–AI collaboration in auditing. Finally, we discuss the potential to embed tools like Vipera into on-the-ground responsible AI workflows in industry settings and beyond. Together, these opportunities outline design directions for building future auditing tools capable of scaling to the complexity and diversity of generative AI systems, while enhancing and complementing auditors’ capacity of sensemaking AI outputs at scale.